Determine the resultant of the following: 9.0 N at 40°, 8.0 N at 120°, and 14.0 N at 22 0°, all




angles being measured with the positive x-axis. Use component method in finding the resultant.

1
Expert's answer
2022-01-26T10:27:46-0500

Determine the resultant of the following: 9.0 N at 40°, 8.0 N at 120°, and 14.0 N at 22 0°, all angles being measured with the positive x-axis. Use component method in finding the resultant.


"F_x=9\cos40°+8\cos120°+14\cos220°=\\=-7.83\text{ N}.\\\space\\
F_y=9\sin40°+8\sin120°+14\sin 220°=\\=3.71\text{ N}.\\\space\\
F_R=\sqrt{F_x^2+F_y^2}
=8.6\text{ N},\\\space\\
\theta=180°+\arctan\dfrac{F_y}{F_x}
=155°"